Find the nature of the planar section of the conicoid x
2
3 −
y
2
4 = z by the plane
x+2y−z = 6
1
Expert's answer
2020-03-05T16:52:35-0500

"\dfrac{x^2}{3}-\dfrac{y^2}{4}=z"

"x+2y\u2212z=6"

We will find z

"z=x+2y\u22126"


"\dfrac{x^2}{3}-\dfrac{y^2}{4}=x+2y−6"

"4x^2-3y^2-12x-24y=-72\\ 4(x^2-3x)-3(y^2+8y)=-72\\"

"4(x-3)^2-3(y+2)^2=-72+9-48=-111\\"


"\dfrac{4(x-\dfrac{3}{2})}{111}-\dfrac{3(y+4)^2}{111}=-1\\ \dfrac{(x-\dfrac{3}{2})}{\dfrac{111}{4}}-\dfrac{(y+4)^2}{{\dfrac{111}{3}}}=-1\\"

This is a equation of conjugated hyperbola

"center 
(\frac{3}{2};-4)"

"semi-axes :
\\a=\sqrt{\frac{111}{4}}, b=\sqrt{\frac{111}{3}}"